OpenStudy (anonymous):

-( )y^2 − [-5y − y(-7y − 9)] − [-y (15y + 4)] = 0 How do I find the missing term?

OpenStudy (anonymous):

Oh gosh, I'm sorry, i have no idea =) @pooja195 @phi

OpenStudy (phi):

you will want to combine like terms.... simplify that mess until you can see what is what.

OpenStudy (phi):

for example, tackle the last term: [-y (15y + 4)] distribute the -y (that means multiply -y times each term inside the parens) can you do that ?

OpenStudy (anonymous):

-15y^2-4y right?

OpenStudy (phi):

yes we have to be careful. so far we have -( )y^2 − [-5y − y(-7y − 9)] − (-15y^2-4y) = 0 I would "distribute the -1" in − (-15y^2-4y) which is the same as -1* (-15y^2-4y) what do we get ?

OpenStudy (phi):

refresh your browser if that did not display correctly

OpenStudy (anonymous):

15y^2+4y ?

OpenStudy (phi):

ok. now we have -( )y^2 − [-5y − y(-7y − 9)] +15y^2 + 4y = 0 now tackle the middle stuff: − [-5y − y(-7y − 9)]

OpenStudy (anonymous):

is it (5y-y)(-7y-9)?

OpenStudy (phi):

if we distribute the -1 out front: −1* [-5y − y(-7y − 9)] = -1* -5y + -1*-y *(-7y-9)

OpenStudy (phi):

no it's not (5y-y)(-7y-9) you should get 5y +y(-7y-9) and now distribute the y

OpenStudy (anonymous):

5y-7y^2-9y ?

OpenStudy (phi):

yes. now it's -() y^2 + 5y-7y^2-9y+15y^2 + 4y = 0 now look for "like terms" which means y to the same power ignoring the first term (for the moment) we have -7y^2 + 15y^2 can you simplify that ?

OpenStudy (anonymous):

8y^2

OpenStudy (phi):

and we have 5y -9y +4y which adds up to 0y or just 0

OpenStudy (phi):

- () y^2 +8y^2 = 0

OpenStudy (anonymous):

9y^2=0

OpenStudy (phi):

I think the () means fill in a number so you get 0 out of the equation in other words, replace () with a number.

OpenStudy (phi):

I would pick 8

OpenStudy (anonymous):

-9y^2 ?

OpenStudy (phi):

because -8y^2 + 8y^2 = 0

OpenStudy (phi):

the idea is we want it to be true no matter what value y is

OpenStudy (anonymous):

o ok . JUst 8 or -8y^2

OpenStudy (phi):

they ask -( )y^2 − [-5y − y(-7y − 9)] − [-y (15y + 4)] = 0 How do I find the missing term? It looks like if we put in 8 where the () is, then we will get 0 out (after simplifying) so just 8

OpenStudy (anonymous):

Much obligied @phi